Factor this equation
9c^3-12c^2+18c-24

Answer by stanbon(75887)

Factor this equation
9c^3-12c^2+18c-24
-------------------
= 3(3c^3-4c^2+6c-8)
----------------------------
factor the 1st 2 and the last 2 terms separately:
= 3(c^2(3c-4)+6(3c-4)
------------------------
= 3(3c-4)(c^2+6)
======================
Cheers,
Stan H.
===========